OpenStudy (anonymous):

Why is the interval of convergence, for the Taylor series about x=0 for f(u)=(1+u)^(-1/2), equal to |u|<1 ?

OpenStudy (anonymous):

\[f(u)=(1+u)^{-\frac{1}{2}}\]

OpenStudy (anonymous):

@satellite73 @experimentX

OpenStudy (experimentx):

expand that binomially ..

OpenStudy (anonymous):

what would the result of that be?

OpenStudy (experimentx):

\[ \frac{1}{{\sqrt {1 + x} }} = \sum\limits_{n = 0}^\infty {\frac{{{{( - 1)}^{n + 1}}}}{{{2^{2n}}}}} {2n\choose n}{x^n} \]

OpenStudy (anonymous):

sorry...not understanding how this is helping

OpenStudy (experimentx):